### v3

# Pavlok-Nudge: A Feedback Mechanism for Atomic Behaviour Modification with Snoring Usecase

Abstract
This paper proposes a feedback mechanism to change behavioural patterns using the Pavlok device. Pavlok utilises beeps, vibration and shocks as a mode of aversion technique to help individuals with behaviour modification. While the device can be useful in certain periodic daily life situations, like alarms and exercise notifications, the device relies on manual operations that limit its usage. To automate behaviour modification, we propose a framework that first detects targeted behaviours through a lightweight deep learning model and subsequently nudges the user through Pavlok. Our proposed solution is implemented and verified in the context of snoring, which captures audio from the environment following a prediction of whether the audio content is a snore or not using a 1D convolutional neural network. Based on the prediction, we use Pavlok to nudge users for preventive measures, such as a change in sleeping posture. We believe that this simple solution can help people to change their atomic habits, which may lead to long-term health benefits. Our proposed real-time, lightweight model (99.8% less parameters over SOTA; 1,278,049 $\rightarrow$ 1337) achieves SOTA performance (test accuracy of 0.99) on a public domain benchmark. The code and model are publicly available at https://github.com/hasan-rakibul/pavlok-nudge-snore.
Index Terms:
Pavlok, behaviour, deep learning, health and wellbeing, snore.
I Introduction
Software-based Behavioural intervention technologies (BITs) have the potential to transform human atomic behaviour [1], which could further lead to long-term health benefits that directly impact human cognition and emotional states [2]. The emerging field of BITs in affective computing mainly includes websites, mobile applications and Internet of Things (IoT) sensor-based designs aiming to make people’s lives better [3, 4]. To this end, we leverage the Pavlok [5], a wearable device aimed to help people ‘break their behavioural pattern’ using aversive conditioning. It accomplishes this by providing nudges in the form of beeps, vibration and/or electric shock to the wearer’s wrist when they engage in targeted behaviours, for example, snoring. The main intention behind this approach is that over time the wearer will become conscious of the atomic behaviour with the unpleasant sensation of the beep/vibration/shock. Even over time, the individual will be less inclined to repeat the behaviour. Moreover, Pavlok can also work in synchronisation with a smartphone application to monitor progress by setting up behaviour change goals. While the native Pavlok is generally useful in certain periodic daily life situations, like alarms, exercise notifications and nail-biting, the device mostly relies on manual operation by the user rather than automatic action detection, which limits its usefulness. Similarly, it lacks the advanced sensory capabilities to detect and take action on relevant events in real time.
Figure 1: Pavlok-Nudge Overview. Pavlok-Nudge is an interactive, nearly real-time software application aiming to aid individuals in atomic behaviour modification. Here, we present the use case for snoring behaviour modification. Our proposed solution consists of the wrist-borne Pavlok and a smartphone (see 1, 2, 3). The microphone on the mobile device captures audio, which is further passed through a deep learning-based framework to detect snoring sounds. Upon snoring detection, the mobile device provides a signal to the Pavlok to nudge the user for posture change to stop snoring (see 2). Scenarios (1), (2) and (3) happen sequentially over time.
Motivation. Snoring is the harsh sound that occurs during sleep when the throat muscles of a person relax, causing narrowness in the airway, which triggers snoring as the relaxed airway vibrates. According to medical research, a natural way to reduce or avert snoring is to sleep on the side [6]. However, it is difficult to maintain the posture involuntarily while a person sleeps. Thus, we believe that our proposed solution, Pavlok-Nudge, can help individuals with behaviour training. Statistically, approximately 75% of people who snore develop obstructive sleep apnea [7, 6]. Sleep apnea leads to the risk of several health issues such as heart disease, diabetes, obesity, arrhythmia (irregular heart rhythm), headache and stroke [8]. Poor quality of sleep reduces memory, thinking capability and proper mental functioning in daily life situations. Prolonged compromises in sleep can lead to frequent irritability, short-temperateness, and even depression [8, 9].
For our use case, the mobile phone’s microphone is used to capture snoring noises. The phone can establish a Bluetooth Low Energy (BLE) connection with the Pavlok device. In this wireless pathway, the phone then sends a signal to the Pavlok device to deliver a ‘nudge’ or ‘shock’ if snoring is detected. Our proposed solution augments the sensory capabilities in the Pavlok. The proposed solution also allows greater customisation and control over Pavlok’s responses related to context. For example, a quiet snore may elicit a vibrate, but a loud snore a shock, and so on. Users can switch between beep, vibration and shock, along with the intensity of the nudge stimulus. Thus, overall, it is user-friendly and adjustable to best suit each individual’s needs. The main contribution of the paper is as follows:
  • Pavlok-Nudge is a software-based behaviour intervention technology to help individuals to break behavioural patterns. It is an interactive, real-time, open-sourced solution that will help the research community to extend it to other human behaviour modification settings.
  • To the best of our knowledge, we are the first to develop a nudge-based approach to snoring-based behaviour modification. On a dataset of 1,000 snoring and non-snoring sounds, our lightweight model (99.8% less parameters over SOTA; 1,278,049 $\rightarrow$ 1,337) achieves a SOTA test accuracy of 0.99.
TABLE I: Where we stand now?: Overview of deep learning algorithms for snoring detection.
Ref. Year Dataset Details Avail. Feature Classifier
[10] 2015 # Sub: 15 $\times$× Raw f-MLP
[11] 2017 # Sub: 68, Lab $\times$× Freq. MLP
[12] 2018 # Sub: 20, Hospital $\times$× MFCC RNN
[13] 2019 # Sub: 10, Bedroom $\times$× Raw 1D-CNN
[14] 2019 # Sam: 1000, Web MFCC 2D-CNN
[15] 2020 # Sub: 15, Hospital $\times$× Mel-spec. CNN-LSTM-DNN
[16] 2021 # Sub: 38, Lab $\times$× CQT-Spec. CNN-LSTM
[17] 2022 # Sam: 1000 Raw NA
[18] 2023 # Sub: 88, Hospital $\times$× Raw 1D/2D-CNN
Our 2024 [17, 14] MFCC 1D-CNN
Figure 2: Pavlok-Nudge Workflow. The Pavlok-Nudge framework senses the environment using the microphone of a mobile or laptop device. To process the audio input, a non-overlapping sliding window-based approach divides the input audio into 1-second chunks, which are then fed to a deep learning-based snore detection model. After collecting 10 such chunks’s output (i.e., 10 seconds of audio), a voting strategy is used to determine if the incoming signal belongs to the snore category. Finally, upon reaching a decision, the system utilises a WiFi signal to connect with the Pavlok device and deliver a user-specific ‘nudge’. For additional details about the workflow, please refer to Section III.
II Related work
Snore Detection and Localisation. Prior literature on snore detection using machine learning methods mainly focuses on detection aspects without any intervention strategy (See Table I for more details). The earliest version of snore recording and detection software was developed by SnoreLab [19]. However, the interface does not provide a snore prevention strategy to act upon snore detection. Non-deep learning-based methods such as logistic regression classifier with auditory image modelling features [20] and AdaBoost classifier with spectral features [21] were utilised for snore detection. Recently, Romero et al. [22] and Emoto et al. [23] use latent space feature from a pre-trained deep learning model to predict if the incoming audio signal is a snore. Similarly, Arsenali et al. [12] use a recurrent neural network with mel-frequency cepstral coefficients (MFCC) features to detect snoring. The most closely related work by Khan [14] also utilises a convolutional neural network (CNN)-based deep learning framework with MFCC feature to detect snoring audio. The proposed solution also includes a vibration device to alert the user to change sleeping posture. However, the deployment situation consists of wired signals, which causes uncomfortable situations for the user. To this end, our solution is more user-friendly and easily deployable.
Pavlok for ‘Breaking Behavioural Pattern’. We leverage the advantage of Pavlok to alert users to change behavioural patterns. The design of the Pavlok device alerts the user periodically or in prior defined time gaps. However, snoring and relevant behaviours require real-time monitoring. For instance, a person who is sleeping cannot pre-define their snoring ahead of time to generate a nudge. To this end, we develop a software solution equipped with the ability to detect and respond to such behavioural events automatically. With the integration of our behaviour-specific deep learning model, the Pavlok is now able to aid individuals in modifying atomic behaviour, particularly which occurs involuntarily.
Figure 3: Receiver operating characteristic curve on the test sets.
III Pavlok-Nudge Pipeline
Our proposed ‘Pavlok-Nudge’ solution consists of a deep learning-based snore-detection model, mobile application and Pavlok (See Figure 2).
III-A Deep Learning-Based Snore Detection Model
Input. Given the continuous sensing of audio using the mobile’s microphone, we use a non-overlapping 1-second audio chunk for our analysis. The deep learning-based framework used in our proposed system takes 1-second audio segments to detect the occurrence of snoring incidents. Following the recommendation from previous work [24], we extract 25 MFCC components. For each 1-second audio segment, we compute MFCCs at 85 evenly spaced time points, resulting in a $25\times 8525 × 85$ matrix representation for each audio sample.
Model Architecture. For snore detection, we employ a 1D CNN architecture. After evaluating various configurations of convolutional and fully connected layers, we finalised a model with two convolutional layers, each using $3\times 33 × 3$ kernels, followed by a fully connected layer. Each convolutional layer is followed by a ReLU activation function and max-pooling. The first convolutional layer outputs 8 channels, while the second expands this to 16 channels, allowing for progressively richer feature extraction from the input audio signal. Max-pooling layers with $2\times 22 × 2$ kernels reduce the spatial dimensions by half. After the convolutional layers, the network output is flattened and passed through a fully connected layer. To mitigate overfitting, a dropout layer with a 0.4 rate is applied before the final classification layer. The output is then passed through a Sigmoid activation function to produce a probability for binary classification. This lightweight architecture effectively captures both temporal and frequency features of snore sounds while maintaining a compact footprint of just 1,337 trainable parameters.
Model Training. The model is trained using the Adam optimiser with a learning rate of 0.001 to minimise the binary cross-entropy loss. Training is capped at a maximum of 100 epochs. We tuned various hyperparameters, including the learning rate, dropout rate and convolutional layer sizes, to optimise the architecture. To prevent overfitting and ensure optimal generalisation, we implement an early stopping criterion, halting the training if the validation loss does not improve for three consecutive epochs. The best-performing model, identified by the lowest validation loss during training, is saved to ensure the most effective version is retained for subsequent inferencing.
Dataset. We utilise two snoring datasets. (1) Khan [14] releases a collection of 500 snoring and 500 non-snoring sounds obtained from various online sources. Following the elimination of silent segments from the audio recordings, individual samples lasting one second each are further generated. The snoring sounds are derived from individuals spanning different age groups, including children, adult men and adult women. Among the 500 snoring samples, 72.6% were characterised by the absence of any background noise, while the remaining 27.4% exhibited non-snoring background noise. As for the 500 non-snoring samples, the dataset has encompassed 50 samples for each of the 10 distinct ambient sounds commonly encountered near individuals during sleep, such as a baby crying, clock ticking, toilet flushing, sirens, television noise, car noise, people talking, rain and thunderstorms. Following [14], we split the dataset into three partitions: 700 samples for training, 200 samples for validation and the remaining 100 samples for testing the model.
(2) MaleFemale [17] is another dataset of 1,000 snoring sounds, including 500 from males and 500 from females, is available at Kaggle. As there are no non-snoring sounds in this dataset, we merged this dataset with Khan [14], resulting in a dataset of 2,000 samples. Like Khan [14], we randomly selected 100 samples from the MaleFemale dataset and made a consolidated test set of 200 samples (i.e., 100 from Khan and 100 from the MaleFemale dataset). This merged dataset includes 1,600 training samples (700 from Khan and 900 from the MaleFemale dataset).
Evaluation Protocol. We report the performance of our model in terms of standard classification metrics, including classification accuracy, precision, recall, F1 score and area under the receiver operating characteristic curve (AUROC).
Results. The performance of our snore detection model across different configurations is presented in Table II. Our final model achieved an accuracy of 0.99 on the test set, significantly outperforming the baseline model proposed by Khan [14], which achieved an accuracy of 0.96.
In addition to its superior performance, our model maintains an exceptionally low parameter footprint of only 1,337 trainable parameters. The baseline 2D CNN model requires a much larger and more complex architecture yet achieves a lower accuracy. In addition to better accuracy, our model delivers high precision (0.98), recall (1.00), F1 score (0.99) and AUROC (0.99).
Software Front-End. We developed a software interface that employs JavaScript and Bluetooth technology to facilitate behaviour modification using the Pavlok 3 device [5].
III-B Real-Time Monitoring
The detection cycle in our proposed system comprises the continuous recording of a 10-second audio clip using the microphone of a device such as a smartphone or a computer. This audio clip is then segmented into ten 1-second intervals, which are processed through a pre-trained neural network on the device (smartphone or computer). The neural network is designed to detect the occurrence of snoring incidents in the audio clip. Specifically, if seven out of ten segments are identified as containing snoring incidents, then the app delivers the pre-selected stimulus to the user via the Pavlok device. The detection cycle is designed to continuously monitor snoring behaviour and provide a prompt and effective aversive stimulus to facilitate behavioural modification.
III-C Comparison with Existing BITs
The work more closely related to the proposed approach is the vibration-based feedback mechanism proposed by Khan [14]. They devised a snoring detection model utilising an MFCC feature-based CNN architecture, attaining a commendable accuracy of 0.96. Firstly, our proposed model outperforms this approach in terms of accuracy. Secondly, our approach involves the integration of a custom-made embedded system affixed to the upper arm of individuals, which beeps or vibrates or shocks upon detecting snoring and continues nudging on an adjustably cycle until the person changes their sleeping position, which we could detect using the Pavlok’s accelerometer. Despite its intended functionality, Khan’s device’s bulkiness and wired connections pose discomfort to the user during sleep, which is not a problem with our system.
TABLE II: Performance of our snore detection model and its comparison with the literature.
Data Model O/p chnl #Params Acc. P R F1 AUC
[14]
2D CNN
[14]
$32-$
64-64
1,278,049 0.96
1D CNN
8-16 1,337 0.99 0.98 1.00 0.99 0.99
4-8 577 0.96 0.93 1.00 0.96 0.96
8 945 0.95 0.94 0.96 0.95 0.95
16 1,873 0.97 0.94 1.00 0.97 0.97
[14]
[17]
1D CNN
8-16 1,337 0.96 0.97 0.98 0.97 0.94
  • O/p chnl: # channels in the output of convolutional layers; P: Precision; R: Recall
IV Discussion and Impact Statement
Impact. While snoring may seem like a minor inconvenience, it can have a significant impact on a person’s health and well-being. One of the most difficult side effects of snoring is sleep disruption, which can lead to daytime fatigue and sleepiness, making it difficult to concentrate and perform daily tasks efficiently. Additionally, snoring can be a sign of sleep apnea, a condition that causes a person to stop breathing periodically throughout the night. This can have serious health implications, including an increased risk of high blood pressure, heart disease and stroke. Snoring can also cause tension and frustration in a sleeping partner, leading to relationship issues and affecting their health by disrupting their sleep. Furthermore, snoring can cause cognitive impairment and memory loss due to a decrease in oxygen levels during sleep. Overall, snoring can have a significant impact on a person’s physical, emotional, and social well-being, and it is important to address the issue promptly to prevent further health complications. Tools such as ours can be useful to assist treatment.
Limitation. While the neural network utilised in our proposed system for snoring detection provides accurate and reliable results, one of the limitations is that it is trained on general snoring clips and cannot differentiate between specific users. As a result, in scenarios where someone is sleeping next to their partner, the Pavlok device might be triggered to deliver a stimulus even if the user is not the one snoring. To tackle the problem, we will integrate speaker diarisation [25] to identify the person from whom the audio is coming in a multi-person scenario. The user interface can be revamped to make it more user-friendly. We do not claim that this method can clinically or medically treat snoring. Instead, we present a tool to assist in the personal management of snoring. The tool will need to be clinically tested before we can make any such claims.
Ethical Use and Privacy Concerns. Users are expected to respect the licenses of the deployed components used in Pavlok-Nudge. During run-time, Pavlok-Nudge detects and tracks audio signals in the user’s living space. This could potentially breach the privacy of the subjects present in the environment. Thus, the user must obtain consent from the subjects before using our application. While developing this solution, we make sure that Pavlok-Nudge does not save any audio in the file system. Moreover, all the operations are performed in real-time; buffered audio signals are removed immediately after prediction.
V Conclusion
To the best of our knowledge, the proposed Pavlok-Nudge workflow is one of the first interactive behaviour intervention strategies for atomic behaviour modification. At its core, a deep learning model detects the targeted behavioural pattern and nudges the user through a Pavlok device. Our experiments on the snoring usecase results in commendable performance – achieving SOTA test performance – using a lightweight 1D CNN architecture of only 1,337 trainable parameters. The proposed software application can be easily transferable across different OS platforms. We believe that Pavlok-Nudge have the potential to significantly impact the future behaviour intervention technologies.
Acknowledgement
This work was supported by resources provided by the Pawsey Supercomputing Research Centre with funding from the Australian Government and Government of Western Australia. We thank Dr Susannah Soon, Senior Lecturer at Curtin University, for her comments on an initial version of this paper.